A ring (whose mass is m, the average
radius r and has a negligible thickness) is
suspended on support A. If the coefficient of friction between the support and the ring
is μ,
a) Find the vertical P force that can slide the ring on the support.
b) What is the value of the angle β when the ring is on the verge of slide?

B/ P
Answer #1

X90 B q po mg Efn = N = (P+mg) cosß. 0 IT, = 0 => t = (1 +mg) sing - (2 here f = MN 3 from egr f = toop Be ton for meg 27 Brt

Please RATE my answer

Thank you